OpenStudy (1645323):

what is the period of 2sin4(theta)???

OpenStudy (anonymous):

period of \(\sin(b\theta)\) is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

in your case \(b=4\)

OpenStudy (anonymous):

y = a sin (bx+c) + k

OpenStudy (1645323):

is it pi/2?